# Cosmological scalar and tensor perturbations with a scalar field: quadratic-order effective energy-momentum tensor

### v4

@@ II.1 Perturbations
where $V_{\phi}=dV/d\phi$.
@@ II.2 Linear solutions
where the overdot represents the derivative with respect to the cosmological time $t$, ˙$\dot{}\equiv d/dt=(1/a)d/d\eta$, and $H=\dot{a}/a$.
@@
$\Psi_{\bf k}(\eta)\approx\left\{\begin{array}[]{ll}A_{1}\epsilon&\quad(\text{for long wavelength})\vspace{0.5em}\\ 4\pi G\dot{\phi}_{0}[c_{1}\sin(k\eta)+c_{2}\cos(k\eta)]&\quad(\text{for short wavelength})\end{array}\right.,$ (20)
where $A_{1}$ and $c_{i}$ are complex constants, and $k=\sqrt{{\bf k}\cdot{\bf k}}$. Note that $\Psi_{\bf k}$ is approximately constant in the long-wavelength limit and rapidly oscillates in the short-wavelength limit. The long-wavelength limit corresponds to $k\ll\cal H$ for which the physical wavelength $\lambda_{\text{phys}}=a/k$ of the perturbation modes is much larger than the Hubble scale $H^{-1}$, $\lambda_{\rm phys}\gg H^{-1}$. In the inflationary stage, the scale factor is approximated by $a(\eta)\approx-1/(H\eta)$, where $\eta<0$. Then one gets ${\cal H}\approx-1/\eta$, and the condition for the long-wavelength limit becomes $k\ll{\cal H}\to-k\eta\ll 1$. The short-wavelength limit corresponds to $k\gg{\cal H}$ which reduces to $-k\eta\gg 1$ in the inflationary stage.
@@
[…]
@@
where $\psi^{(n)}(z)=d^{n}\psi(z)/dz^{n}=d^{n+1}\log\Gamma(z)/dz^{n+1}$. The solution has been expanded in the small quantities such as $\epsilon$ and $k\eta=-\sigma_{L}$.
@@
where ${\cal SC}\equiv b_{1}\sin(k\eta)+b_{2}\cos(k\eta)$ and ${\cal CS}\equiv b_{1}\cos(k\eta)-b_{2}\sin(k\eta)$. The solution has been expanded in the small quantities such as $\epsilon$ and $(k\eta)^{-1}=-\sigma_{S}$.
@@ III Second-order effective energy-momentum tensor (2EMT)
[…]
@@ III.2.1 Long-wavelength limit
$\hat{\tau}_{00}$ $=-\frac{224G|b_{2}|^{2}}{k\eta^{2}a^{2}}-\frac{256G{\rm Re}(b_{1}b_{2}^{*})}{\eta a^{2}}-\frac{256G|b_{2}|^{2}\epsilon^{2}}{3k^{3}\eta^{4}a^{2}}+\frac{64G|b_{2}|^{2}\epsilon}{k\eta^{2}a^{2}}\left[7\log(-k\eta)+4-7\log 2-7\psi^{(0)}(3/2)\right]$ (37)
@@ III.3 Scalar-Tensor coupled ($\cal{ST}$)
$\hat{\tau}_{ij}$ $=h^{\prime}_{ij}\left(Q^{\prime\prime}+2\mathcal{H}Q^{\prime}+2\mathcal{H}^{\prime}Q+\Delta Q-2\Delta E^{\prime}\right)$ (45)
$+h_{ij}\left\{4\left(\mathcal{H}+a^{2}\frac{V_{\phi}}{\phi_{0}^{\prime}}\right)\Psi^{\prime}+4\left(\mathcal{H}^{\prime}+2\mathcal{H}^{2}+a^{2}\mathcal{H}\frac{V_{\phi}}{\phi_{0}^{\prime}}\right)\Psi\right.$ (46)
$+2\mathcal{H}Q^{\prime\prime}+4\left(\mathcal{H}^{\prime}+\mathcal{H}^{2}\right)Q^{\prime}+2\Delta Q^{\prime}+2\mathcal{H}\Delta Q-2\left[3\mathcal{H}\mathcal{H}^{\prime}+2\mathcal{H}^{3}-4\pi Ga^{2}(3\mathcal{H}V_{0}+2\phi_{0}^{\prime}V_{\phi})\right]Q$ (47)
@@
For 2EMT of $\cal{ST}$ terms, the nonzero components are $\hat{\tau}_{11}=-\hat{\tau}_{22}\equiv\hat{\tau}_{\rm ST}$ for $+$-polarization, and $\hat{\tau}_{12}=\hat{\tau}_{21}=\hat{\tau}_{\rm ST}$ for $\times$-polarization of the tensor mode.
@@
[…]
@@ IV Conclusions
Future research could include extending this framework to exploring the scalar-tensor interaction in alternative gauges, and assessing observational signatures of 2EMT contributions, such as their imprints on the cosmic microwave background and gravitational wave spectra. Involving with the tensor mode, there could be a quantum effect [15] which was absent in the FRW universe with only the scalar-mode perturbation [1], which is worthwhile to investigate.

### v5

@@ I Introduction
This is regarded as the source of the back-reaction to the geometry part $G_{\mu\nu}^{(1)}[g^{(2)}]$ in Eq. (1), while the linear-order term $T_{\mu\nu}^{(1)}[g^{(2)},\delta\phi^{(2)}]$ has stochastically vanishing effect when it is integrated in time [3, 4].
@@ II.1 Perturbations
where $V_{\phi}=dV/d\phi$. Here, we used the relations obtained from the first-order equations for the gauge-invariant variables, $\Psi=\Phi$, and
$\overline{\delta\phi}=\frac{\Psi^{\prime}+\mathcal{H}\Psi}{4\pi G\phi_{0}^{\prime}},\quad\overline{\delta\phi}^{\prime}=\frac{\Delta\Psi-K\Psi^{\prime}-L\Psi}{4\pi G\phi_{0}^{\prime}}.$ (13)
@@ II.2 Linear solutions
where the overdot represents the derivative with respect to the cosmological time $t$, $\dot{}\equiv d/dt=(1/a)d/d\eta$, and $H=\dot{a}/a$.
@@
$\Psi_{\bf k}(\eta)\approx\left\{\begin{array}[]{ll}A_{1}\epsilon&\quad(\text{for long wavelength})\vskip 5.0pt\\ 4\pi G\dot{\phi}_{0}[c_{1}\sin(k\eta)+c_{2}\cos(k\eta)]&\quad(\text{for short wavelength})\end{array}\right$., (21)
where $A_{1}$ and $c_{i}$ are complex constants, and $k=\sqrt{{\bf k}\cdot{\bf k}}$. Note that $\Psi_{\bf k}$ is approximately constant in the long-wavelength limit and rapidly oscillates in the short-wavelength limit. The long-wavelength limit corresponds to $k\ll\cal$ H for which the physical wavelength $\lambda_{\text{phys}}=a/k$ of the perturbation modes is much larger than the Hubble scale $H^{-1}$, $\lambda_{\rm phys}\gg H^{-1}$. In the inflationary stage, the scale factor is approximated by $a(\eta)\approx-1/(H\eta)$, where $\eta<0$. Then one gets ${\cal H}\approx-1/\eta$, and the condition for the long-wavelength limit becomes $k\ll{\cal H}\to-k\eta\ll 1$. The short-wavelength limit corresponds to $k\gg{\cal H}$ which reduces to $k\eta\gg 1$ in the inflationary stage.
@@
$\sigma_{L}$ $\equiv-k\eta\ll 1\quad\mbox{(for long wavelength)},$ (25)
$\sigma_{S}$ $\equiv-(k\eta)^{-1}\ll 1\quad\mbox{(for short wavelength)}.$ (26)
@@
where $\psi^{(n)}(z)=d^{n}\psi(z)/dz^{n}=d^{n+1}\log\Gamma(z)/dz^{n+1}$. The solution has been expanded in the small quantities such as the slow-roll parameter $\epsilon$ and the wave-length parameter $k\eta=-\sigma_{L}$ in the order of ${\cal O}(\varepsilon^{-1\sim 2})$ where $\varepsilon=\epsilon$, or $\sigma_{L}$. Afterwards the expression has been recovered, $\sigma_{L}\to-k\eta$.
@@
where ${\cal SC}\equiv b_{1}\sin(k\eta)+b_{2}\cos(k\eta)$ and ${\cal CS}\equiv b_{1}\cos(k\eta)-b_{2}\sin(k\eta)$. The solution has been expanded in the small quantities $\epsilon$ and the wave-length parameter $(k\eta)^{-1}=-\sigma_{S}$ in the order of ${\cal O}(\varepsilon^{0\sim 4})$ where $\varepsilon=\epsilon$, or $\sigma_{S}$. Afterwards the expression has been recovered, $\sigma_{S}\to-(k\eta)^{-1}$.
@@ III Second-order effective energy-momentum tensor (2EMT)
where $\tau_{\mu\nu}({\bf k})$ means $\tau_{\mu\nu}$ for a given $\bf k$} mode, and $\left\langle\right\rangle$ denotes the “Fourier-transformed” quantity with the Fourier modes introduced in Eqs. (15) and  (16). We expand $\hat{\tau}_{\mu\nu}$ in terms of the small parameter $\varepsilon$ such as the slow-roll parameters $\epsilon$ and $\delta$, and the wave-length parameters $\sigma_{L}$ and $\sigma_{S}$.
@@ III.2.1 Long-wavelength limit
$\hat{\tau}_{00}$ $=-\frac{224G|b_{2}|^{2}}{k\eta^{2}a^{2}}-\frac{256G{\rm Re}(b_{1}b_{2}^{*})}{\eta a^{2}}-\frac{256G|b_{2}|^{2}\epsilon^{2}}{3k^{3}\eta^{4}a^{2}}+\frac{64G|b_{2}|^{2}\epsilon}{k\eta^{2}a^{2}}\left[7\log(-k\eta)+4-7\log 2-7\psi^{(0)}(3/2)\right]$
@@ III.3 Scalar-Tensor coupled ($\cal{ST}$)
$\hat{\tau}_{ij}$ $=h^{\prime}_{ij}\left(Q^{\prime\prime}+2\mathcal{H}Q^{\prime}+2\mathcal{H}^{\prime}Q+\Delta Q-2\Delta E^{\prime}\right)$
$+h_{ij}\left\{4\left(\mathcal{H}+a^{2}\frac{V_{\phi}}{\phi_{0}^{\prime}}\right)\Psi^{\prime}+4\left(\mathcal{H}^{\prime}+2\mathcal{H}^{2}+a^{2}\mathcal{H}\frac{V_{\phi}}{\phi_{0}^{\prime}}\right)\Psi\right$.
$+2\mathcal{H}Q^{\prime\prime}+4\left(\mathcal{H}^{\prime}+\mathcal{H}^{2}\right)Q^{\prime}+2\Delta Q^{\prime}+2\mathcal{H}\Delta Q-2\left[3\mathcal{H}\mathcal{H}^{\prime}+2\mathcal{H}^{3}-4\pi Ga^{2}(3\mathcal{H}V_{0}+2\phi_{0}^{\prime}V_{\phi})\right]Q$
@@
For 2EMT of $\cal{ST}$ terms, the nonzero components are $\hat{\tau}_{11}=-\hat{\tau}_{22}\equiv\hat{\tau}_{\rm ST}$ for ++-polarization, and $\hat{\tau}_{12}=\hat{\tau}_{21}=\hat{\tau}_{\rm ST}$ for $\times$-polarization of the tensor mode. The contributions to $\hat{\tau}_{ij}$ are from the geometric part $G_{\mu\nu}^{(2)}[g^{(1)}]$ and the matter part $T_{\mu\nu}^{(2)}[g^{(1)},\delta\phi^{(1)}]$ in Eq. (2). The matter-part contribution is given by
$\hat{\tau}_{ij}\left[T_{\mu\nu}^{(2)}\right]=h_{ij}\left\{4\Delta\Psi-4\left(3\mathcal{H}+2a^{2}\frac{V_{\phi}}{\phi_{0}^{\prime}}\right)\Psi^{\prime}-4\left(3\mathcal{H}^{2}+2a^{2}\mathcal{H}\frac{V_{\phi}}{\phi_{0}^{\prime}}\right)\Psi+4\left(3\mathcal{H}\mathcal{H}^{\prime}-3\mathcal{H}^{3}-8\pi Ga^{2}\phi_{0}^{\prime}V_{\phi})\right)Q\right\}.$ (45)
@@
(ii) spatially-flat gauge: $\psi=E=0$ $\quad\Rightarrow\quad$ $Q=\Psi/{\cal H}$
(iii) comoving gauge: $\delta\phi=E=0$ $\quad\Rightarrow\quad$ $\Psi=\psi+{\cal H}\beta$ and $Q=(\Psi^{\prime}+{\cal H}\Psi)/({\cal H}^{\prime}-{\cal H}^{2})$
@@ IV Conclusions
Future research could include extending this framework to exploring the scalar-tensor interaction in alternative gauges, and assessing observational signatures of 2EMT contributions, such as their imprints on the cosmic microwave background and gravitational wave spectra. At the current stage, however, it is not very plausible to discuss the observational implications because there still exist gauge issues; 2EMT depends on the gauge choices. Studying more gauge conditions might help in understanding the gauge issue in a relation to the observational implications.
It is also worthwhile to investigate the quantum effect in the future. Involving with the tensor mode, there could be a quantum effect [15] which was absent in the FRW universe with only the scalar-mode perturbation [1].
Our investigation can be extended to the post-inflationary (preheating) stage at which the scalar field exhibits the oscillatory behaviour. A massive-scalar field ($V=m^{2}\phi^{2}/2$), for example, behaves as a pressureless dust in the oscillatory regime, and the scale factor becomes $a(\eta)\propto\eta^{2}$. The set-up will then be that of the perturbation of a scalar field in the Friedmann universe. This will be a mixture of our previous [1, 2] and current work. The perturbations produced at this stage have very short wavelengths, and come back into the horizon at the early stage of the Friedmann universe. The curvature perturbation does not grow much, and its effect is not significant to leave a signature in the density perturbation.
When the second scalar field is introduced, however, the situation may become interesting. When another massive-scalar field is introduced, it will also play as a pressureless dust, and the background universe will still settle down to a Friedmann universe. When two fields are decoupled, they will not produce significant effects. When two fields are coupled, however, the parametric resonance may induce a nontrivial effect such as the well-known “nongaussianity” [16]. Regarding that the nongaussianity is a high-order effect, this set-up might induce an effect also in 2EMT that we consider. It will be interesting to investigate 2EMT in this set-up in the future.
In principle, adding another scalar field to our set-up is straightforward. However, the gauge invariant variables $\overline{\delta\phi_{i}}$ cannot be solved in a simple form by the Bardeen variable as was done for the single field in Eq. (13). Therefore, it is not easy at once to expect what new effect will be brought to us by the second field, in particular, when we consider a second-order effect such as 2EMT. However, it will be a nice attempt to extended our work to the preheating stage with multiple-scalar fields.